niyhawhite:

Which decimal number best describes the fraction fraction 137 over 6

CancerGirl:

Hope this helps u!!!!!

jhonyy9:

\[\frac{ 137 }{ 6} = ? \] do you know what mean decimal number ?

jhonyy9:

@CancerGirl ATTENTION please in the first line you ve wrote 28,333333 and this not is true -please correct it thank you in advance

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @CancerGirl 137/6=137รท6=28.3333... the decimal is being repeated over and over again, so it won't stop. So, the answer is a repeating decimal 22.83 \(\color{#0cbb34}{\text{End of Quote}}\) here you ve wrote =28,33333...

CancerGirl:

It will be 22.83 with the 3repeating The decimal is being repeated over and over again, so it won't stop. So, the answer is a repeating decimal 22.83

CancerGirl:

Hope this helps you!!!

jhonyy9:

let you know that in this case of decimal numbers when a number repeating so this noted with parentheses in this way 137/6 = 22,8(3)

CancerGirl:

Okay